Sandy is 11 kilometers away from Gary. Both begin to walk toward each other at the same time. sandy walks at 4 kilometers per hour. They will meet in 2 hours. How fast is Gary walking?
I know this is a linear equation. I am just not sure how to put it together.

Answer by Fombitz(13823)

Rate*Time=Distance
S*t+G*t=11
4*2+G*2=11
8+2G=11
2G=3
G=3/2 kph